OpenStudy (anonymous):

-42s^4 t^6+36s^3 t^5+6s^2 t^4+24st^2/2st^2 Find the quotient

terenzreignz (terenzreignz):

\[\Large \frac{-42s^4 t^6+36s^3 t^5+6s^2 t^4+24st^2}{2st^2}\] Okay, this is one of the more basic skills in algebra, so pay attention :) First thing you want to do is deal with the numbers, if you can... You'll want to divide each of these coefficients \[\Large \frac{\color{blue}{-42}s^4 t^6+\color{blue}{36}s^3 t^5+\color{blue}{6}s^2 t^4+\color{blue}{24}st^2}{2st^2}\] By the coefficient in the denominator... \[\Large \frac{-42s^4 t^6+36s^3 t^5+6s^2 t^4+24st^2}{\color{red}2st^2}\]

OpenStudy (anonymous):

but it seems like the awnser would be a small simple one. and not that..? is there any way to like simlify that or something??

terenzreignz (terenzreignz):

We are simplifying it... I'm just showing you the steps... one at a time :)

terenzreignz (terenzreignz):

You there? Divide those coefficients at the numerator (the blue ones) by 2 (the coefficient in the denominator).... What do you get?

OpenStudy (anonymous):

add them all together then divide??

terenzreignz (terenzreignz):

No, just divide... okay, I'll do a simple example, aye? :) \[\Large \frac{25x^3y^2 -10x^2y^3}{5xy}\] First thing I'll do is divide the coefficients at the top \[\Large \frac{\color{blue}{25}x^3y^2 -\color{blue}{10}x^2y^3}{5xy}\] by 5...\[\Large \frac{25x^3y^2 -10x^2y^3}{\color{red}5xy}\] we get\[\Large \frac{\color{blue}5x^3y^2 -\color{blue}2x^2y^3}{xy}\] (note how the 5 in the denominator is now gone...) Next, we divide these x's \[\Large \frac{5\color{blue}{x^3}y^2 -2\color{blue}{x^2}y^3}{xy}\]by the lone x in the denominator.... note that to divide powers of x, you simply subtract their exponents... and since the x in the denominator has no exponent, it's understood to have an exponent of 1. So we subtract 1 from all the exponents of x in the numerator... \[\Large \frac{5\color{blue}{x^2}y^2 -2\color{blue}{x}y^3}{y}\](Note how the x in the denominator is gone now, too...) Finally, we divide all the y's in the numerator \[\Large \frac{5{x^3}\color{blue}{y^2} -2{x^2}\color{blue}{y^3}}{y}\] following what we did with the x's, we subtract

OpenStudy (anonymous):

still confused... soooo..? whats the awnser?